# Beyond Pointwise Scores: Decomposed Criteria-Based Evaluation of LLM Responses

### v1

@@ Abstract
Beyond Pointwise Scores:
Decomposed Criteria-Based Evaluation of LLM Responses
@@ 1 Introduction
As large language models (LLMs) are increasingly deployed in high-stakes, expert settings, such as law, medicine, and finance; their outputs must satisfy demanding requirements: factual accuracy, citation support, and coverage of domain-specific obligations Lai et al. (2024a); Wang et al. (2024); Zhang et al. (2024a). However, evaluating such complex, long-form responses remains a fundamental challenge Liang et al. (2023); Chang et al. (2024). Evaluation failures in these domains carry real consequences, including tangible harm, legal liability and erosion of trust in AI systems. To anchor ideas, consider legal question answering, where a lawyer might ask: “Does a \$2B acquisition of a competitor trigger antitrust filing obligations in California?” An effective LLM answer must synthesize statutes, regulations, and case law. Subsequent evaluation is hence a fundamentally open-ended, multi-dimensional challenge.
@@
Unfortunately, existing evaluation paradigms fail to address this challenge. Human expert reviews, though considered the gold standard, are costly and unscalable (e.g., it can take a legal expert around 45 minutes to evaluate one model response by thoroughly checking citations, verifying legal reasoning, and assessing applicability). This approach becomes prohibitively expensive for large-scale evaluation across multiple models and tasks. Lexical metrics like ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002) are cheap but correlate poorly with human judgment on complex, knowledge-intensive tasks. LLM-as-a-judge has emerged as a promising alternative (Zheng et al., 2023; Li et al., 2025; Gu et al., 2025), yet many implementations reduce evaluation to a single pointwise score, obscuring actionable insights. Recent multidimensional variants (e.g., GPTScore Fu et al. (2024a) , G-Eval Liu et al. (2023a)) add labeled axes (e.g., accuracy, completeness) and show improved alignment in general domains; however, they typically rely on generic, task-agnostic criteria that miss domain-specific obligations and hierarchies central to expert settings like law. Checklist-based approaches (e.g., LLM-Rubric (Hashemi et al., 2024), CheckEval (Lee et al., 2024)) increase granularity but require manual rubrics, taxonomies, or seed questions. Hence, while these methods provide greater granularity than naive LLM-as-a-judge, they require substantial manual effort or lack systematic decomposition into dimensions such as precision (factual accuracy, relevance) and recall (coverage), which are essential in expert domains to diagnose specific model behaviors. RAGCHECKER Ru et al. (2025) advances claim-level precision/recall via bidirectional entailment over extracted claims but treats claims uniformly, overlooking requirement hierarchies. Beyond claim overlap, AQuAECHR Weidinger et al. (2025) and ALCE Gao et al. (2023) measure precision and recall of citations to assess citation accuracy and coverage in legal contexts, but are limited to citation verification. A comparison of our work vs related works is illustrated in Table 1.
@@ 3 Experiments
We evaluate DeCE across multiple dimensions: alignment with expert judgment, model-specific diagnostic insights, and reliability of criteria extraction. We instantiate DeCE on legal question answering, a high-stakes domain where LLMs increasingly assist professionals Lai et al. (2024b). DeCE is applicable to other domains with structured, expert-authored outputs (clinical QA, financial compliance, scientific summarization, etc.).
@@
LLMs evaluated. We assess five LLMs generating legal answers under standardized retrieval. These include: general-purpose GPT-4o (OpenAI); reasoning models Gemini-2.5-Pro Comanici et al. (2025) and DeepSeek-R1 DeepSeek-AI et al. (2025); open-source model Llama-3.1-405B Grattafiori et al. (2024); and domain-specific model, Legal Llama-3.1-70B, which we fine-tuned using supervised learning Zhang et al. (2024b) and Direct Preference Optimization Rafailov et al. (2023) on legal corpora. All models have over 70 billion parameters [^2] GPT-4o, DeepSeek-R1, and Llama-3.1-405B were accessed via AWS Bedrock API; Gemini-2.5-Pro via Google Vertex AI API, in accordance with each provider’s terms of service and consistent with their intended use. We fine-tuned Llama-3.1-70B under Meta’s Llama 3 Community License. The finetuning specifics of the Legal Llama-3.1-70B model are out of scope for this paper..
Evaluation method baselines. We compare DeCE (using Claude 3.5 Sonnet Anthropic (2024) as the backbone LLM) against four categories of evaluation methods: (1) Lexical overlap metrics. ROUGE-L and BLEU: standard overlap metrics that are computationally cheap but correlate weakly with semantic correctness on complex generation tasks. (2) Pointwise LLM-as-a-judge. Following prior work (Zheng et al., 2023; Li et al., 2025), we prompt an LLM (Claude 3.5 Sonnet) to assign 0–4 Likert scores relative to the gold answer using a rubric-driven prompt with chain-of-thought Wei et al. (2023); scores are normalized to a GPA-style 0–4 metric (see prompt in Appendix  C.3). (3) Multidimensional LLM-as-a-judge. GPTScore Fu et al. (2024b) and G-Eval Liu et al. (2023a) serve as representative multi-axis judges. For comparability, we: (i) use the same backbone LLM (Claude 3.5 Sonnet), (ii) provide the gold answer as reference , and (iii) report the same dimensions—precision and recall (as in DeCE). (4) Claim-level. RAGCHECKER Ru et al. (2025) which computes claim-level precision/recall via bidirectional entailment. We use author-recommended settings and provide the same gold answers and retrieved materials for consistency. We use Claude 3.5 Sonnet for claim extraction for a fair comparison with DeCE.
@@ B.3 Taxonomic Categories
The legal queries have been systematically categorized by Attorney Editors according to a comprehensive taxonomy capturing the breadth of legal information needs. Our classification framework encompasses various distinct categories organized hierarchically based on query complexity and subject matter. This taxonomic structure enables fine-grained analysis of model performance across different question types and complexity levels.

### v2

@@ 1 Introduction
As large language models (LLMs) are increasingly deployed in high-stakes, expert settings, such as law, medicine, and finance; their outputs must satisfy demanding requirements: factual accuracy, citation support, and coverage of domain-specific obligations (Lai et al., 2024a; Wang et al., 2024; Zhang et al., 2024a). However, evaluating such complex, long-form responses remains a fundamental challenge (Liang et al., 2023; Chang et al., 2024). Evaluation failures in these domains carry real consequences, including tangible harm, legal liability and erosion of trust in AI systems. To anchor ideas, consider legal question answering, where a lawyer might ask: “Does a \$2B acquisition of a competitor trigger antitrust filing obligations in California?” An effective LLM answer must synthesize statutes, regulations, and case law. Subsequent evaluation is hence a fundamentally open-ended, multi-dimensional challenge.
@@
Unfortunately, existing evaluation paradigms fail to address this challenge. Human expert reviews, though considered the gold standard, are costly and unscalable (e.g., it can take a legal expert around 45 minutes to evaluate one model response by thoroughly checking citations, verifying legal reasoning, and assessing applicability). This approach becomes prohibitively expensive for large-scale evaluation across multiple models and tasks. Lexical metrics like ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002) are cheap but correlate poorly with human judgment on complex, knowledge-intensive tasks. LLM-as-a-judge has emerged as a promising alternative (Zheng et al., 2023; Li et al., 2025; Gu et al., 2025), yet many implementations reduce evaluation to a single pointwise score, obscuring actionable insights. Recent multidimensional variants (e.g., GPTScore (Fu et al., 2024a) , G-Eval (Liu et al., 2023a)) add labeled axes (e.g., accuracy, completeness) and show improved alignment in general domains; however, they typically rely on generic, task-agnostic criteria that miss domain-specific obligations and hierarchies central to expert settings like law. Checklist-based approaches (e.g., LLM-Rubric (Hashemi et al., 2024), CheckEval (Lee et al., 2024)) increase granularity but require manual rubrics, taxonomies, or seed questions. Hence, while these methods provide greater granularity than naive LLM-as-a-judge, they require substantial manual effort or lack systematic decomposition into dimensions such as precision (factual accuracy, relevance) and recall (coverage), which are essential in expert domains to diagnose specific model behaviors. RAGCHECKER (Ru et al., 2025) advances claim-level precision/recall via bidirectional entailment over extracted claims but treats claims uniformly, overlooking requirement hierarchies. Beyond claim overlap, AQuAECHR (Weidinger et al., 2025) and ALCE (Gao et al., 2023) measure precision and recall of citations to assess citation accuracy and coverage in legal contexts, but are limited to citation verification. A comparison of our work vs related works is illustrated in Table 1.
@@ 3 Experiments
We evaluate DeCE across multiple dimensions: alignment with expert judgment, model-specific diagnostic insights, and reliability of criteria extraction. We instantiate DeCE on legal question answering, a high-stakes domain where LLMs increasingly assist professionals (Lai et al., 2024b). DeCE is applicable to other domains with structured, expert-authored outputs (clinical QA, financial compliance, scientific summarization, etc.).
@@
LLMs evaluated. We assess five LLMs generating legal answers under standardized retrieval. These include: general-purpose GPT-4o (OpenAI); reasoning models Gemini-2.5-Pro (Comanici et al., 2025) and DeepSeek-R1 (DeepSeek-AI et al., 2025); open-source model Llama-3.1-405B (Grattafiori et al., 2024); and domain-specific model, Legal Llama-3.1-70B, which we fine-tuned using supervised learning (Zhang et al., 2024b) and Direct Preference Optimization (Rafailov et al., 2023) on legal corpora. All models have over 70 billion parameters [^2] GPT-4o, DeepSeek-R1, and Llama-3.1-405B were accessed via AWS Bedrock API; Gemini-2.5-Pro via Google Vertex AI API, in accordance with each provider’s terms of service and consistent with their intended use. We fine-tuned Llama-3.1-70B under Meta’s Llama 3 Community License. The finetuning specifics of the Legal Llama-3.1-70B model are out of scope for this paper..
Evaluation method baselines. We compare DeCE (using Claude 3.5 Sonnet (Anthropic, 2024) as the backbone LLM) against four categories of evaluation methods: (1) Lexical overlap metrics. ROUGE-L and BLEU: standard overlap metrics that are computationally cheap but correlate weakly with semantic correctness on complex generation tasks. (2) Pointwise LLM-as-a-judge. Following prior work (Zheng et al., 2023; Li et al., 2025), we prompt an LLM (Claude 3.5 Sonnet) to assign 0–4 Likert scores relative to the gold answer using a rubric-driven prompt with chain-of-thought (Wei et al., 2023); scores are normalized to a GPA-style 0–4 metric (see prompt in Appendix C.3). (3) Multidimensional LLM-as-a-judge. GPTScore (Fu et al., 2024b) and G-Eval (Liu et al., 2023a) serve as representative multi-axis judges. For comparability, we: (i) use the same backbone LLM (Claude 3.5 Sonnet), (ii) provide the gold answer as reference , and (iii) report the same dimensions—precision and recall (as in DeCE). (4) Claim-level. RAGCHECKER (Ru et al., 2025) which computes claim-level precision/recall via bidirectional entailment. We use author-recommended settings and provide the same gold answers and retrieved materials for consistency. We use Claude 3.5 Sonnet for claim extraction for a fair comparison with DeCE.
@@ B.3 Taxonomic Categories
The legal queries have been systematically categorized by Attorney Editors according to a comprehensive taxonomy capturing the breadth of legal information needs. Our classification framework encompasses various
distinct categories organized hierarchically based on query complexity and subject matter.
This taxonomic structure enables fine-grained analysis of model performance across different question types and complexity levels.